OpenStudy (anonymous):

solve the given system of equation(matrices) x +3y= 6 X-3y=12

OpenStudy (anonymous):

Do you know how to put it in matrix form?

OpenStudy (anonymous):

1 3 x 6 1 -3 y 12 ?

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 3 & x \\ 1 & -3 & y\end{matrix}\right]=\left(\begin{matrix}6 \\ 12\end{matrix}\right)\]

OpenStudy (anonymous):

Now find the inverse of the 2x2 matrix by ignoring the x and y

OpenStudy (anonymous):